# DeepTV: A neural network approach for total variation minimization

### v1

@@ 2. $d=2$
    : Here we will consider two different discretised gradient operators. First the standard forward difference discretized gradient operator denoted by $\nabla^{h}_{\text{F}}$ and the above suggested forward-backward discretisation $\nabla^{h}_{\text{FB}}$. For both choices we incorporate in a standard way respective boundary conditions. The two-dimensional discrete total variations read then as
@@ 4.4 Identifying and solving a numerical issue
We recall that $\mathcal{H}\subset W^{1,1}(\Omega)$ and hence for any $u\in\mathcal{H}$ we have $\operatorname{TV}(u)=\int_{\Omega}|\nabla u|\text{d}x$. We analyze now the $\Gamma$-convergence of (12) with $\operatorname{TV}(u)=\int_{\Omega}\varphi_{\gamma}(\nabla u)\text{d}x$, where $\varphi_{\gamma}(\cdot)\in\left\{|\cdot|_{2,\gamma},\max\{|\cdot|_{2},\gamma\},\phi_{2,\gamma}(\cdot)\right\}$ and how their minimizers behave for $\gamma\to 0$.
@@ 4.4.1 $\Gamma$-convergence of total variation approximations
Let $\varphi_{\gamma}(\cdot)\in\left\{|\cdot|_{2,\gamma},\max\{|\cdot|_{2},\gamma\},\phi_{2,\gamma}(\cdot)\right\}$, $(\gamma_{m})_{m\in\mathbb{N}}>0$ be monotonically decreasing sequences with $\lim_{m\to\infty}\gamma_{m}=0$ and $A\in\{\mathcal{A}_{h}\cap\mathcal{H}_{c},\mathcal{H}_{c}\}$, where $c\geq 0$, $h>0$, and $\mathcal{A}_{h}$ defined as in (20). Further we define the functionals $\mathcal{E}_{m}^{c},\mathcal{E}^{c}:L^{1}(\Omega)\to\overline{\mathbb{R}}$ as
@@
where $A_{n}\in\{\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}},\mathcal{H}_{c_{n}}\}$. Then there exists a sequence $(c_{n})_{n}\subset\mathbb{R}$ such that the sequence $(\mathcal{E}_{n,m})_{n,m}$ $\Gamma$-converge to $\mathcal{E}$ with respect to the $L^{1}(\Omega)$ topology.
@@
$\begin{split}\min_{u_{\theta}\in\mathcal{H}_{c}}\frac{\alpha_{1}}{N}\sum_{x\in\Omega^{h}}|(T^{h}u_{\theta}(x)-g(x)|&+\frac{\alpha_{2}}{N}\sum_{x\in\Omega^{h}}|(T^{h}u_{\theta}(x)-g(x)|^{2}\\ &+\frac{\lambda}{N}\sum_{x\in\Omega^{h}}\widetilde{\varphi}_{\gamma}(\nabla_{\text{FB}}^{h}u(x))\end{split}$ (32)
@@ 4.5 Connection to finite difference discretisation
where $g_{\text{FD}}\in\mathbb{R}^{N}$ is a finite difference discretisation of $g$ such that $g_{\text{FD}}(x)=g(x)$ for all $x\in\Omega^{h}$. In particular, if $\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}=\mathcal{A}_{h_{n}}$ in (24), then (24) is also the integral counterpart of (33). Note that here as in (19) $\nabla^{h}$ and $|\cdot|_{\blacktriangleright}$ might be any suitable discrete gradient and discrete norm. For example, one may choose $\nabla_{\text{F}}^{h}$ and $|\cdot|_{2}$ or $\nabla_{\text{FB}}^{h}$ and $|\cdot|_{2,1}$. Note that if $u_{\text{FD}}(x)=u_{\theta}(x)$ for all $x\in\Omega^{h}$, then $E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$. This observation leads to the following result:
@@
Given the set of points $\{x_{i}\}_{i=1}^{N}=\Omega^{h}\subset\Omega$. If $u_{\text{FD}}\in\mathbb{R}^{N}$ is a solution of (33), then there exists a constant $c\geq 0$ such that $u_{\theta}\in\mathcal{H}_{c}$ is a solution of (19), whereby ${u}_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$.
@@
We construct a mesh $\mathcal{T}$ of simplices with vertices $\{x_{i}\}_{i=1}^{N}$ and let $u_{\text{FD}}$ be a solution of (33). Then there exists a $c\geq 0$ such that $u_{\theta}\in\mathcal{H}_{c}$ is its continuous piecewise affine interpolation on $\mathcal{T}$, i.e. $u_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$. Hence we have $E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$.
Assume that $u_{\theta}$ is not a minimizer of $E_{\theta}$, i.e. there is a $\tilde{u}_{\theta}\in\mathcal{H}_{c}$ with $\tilde{u}_{\theta}\neq{u}_{\theta}$ such that $E_{\theta}(\tilde{u}_{\theta})<E_{\theta}(u_{\theta})$. We define $\tilde{u}_{\text{FD}}\in\mathbb{R}^{N}$ as $\tilde{u}_{\text{FD}}(x_{i})=\tilde{u}_{\theta}(x_{i})$ for $i=1,\ldots,N$. Hence we obtain $E_{\text{FD}}(\tilde{u}_{\text{FD}})=E_{\theta}(\tilde{u}_{\theta})<E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$, which is a contradiction showing that $u_{\theta}$ is indeed a minimizer of $E_{\theta}$. ∎
Note that for any $c\geq 0$ (19) has a solution, see theorem 3.4, while this is not the case for (33). This is why the other direction in proposition 4.7 does not hold in general. Moreover, although [7, Theorem 2.1] states that any continuous piecewise affine function can be represented by a ReLU-NN, it seems not possible to preset a $c\geq 0$ such that $\mathcal{H}_{c}$ contains any continuous piecewise affine functions, not even on a given grid. However, if (33) has a solution then there is a $c\geq 0$, such that the sets of solutions of (33) and (19), respectively, are equivalent in the sense of proposition 4.7.
If $c=\infty$, i.e. the constraint $|\theta|_{\infty}<c$ in (19) is dropped, then we even have the following statement.
@@
Given the set of points $\{x_{i}\}_{i=1}^{N}=\Omega^{h}\subset\Omega$ and let $c=\infty$ in (19), then $u_{\text{FD}}\in\mathbb{R}^{N}$ is a solution of (33), if and only if $u_{\theta}\in\mathcal{H}$ is a solution of (19), whereby ${u}_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$.
@@
$\Rightarrow$”: This direction is shown as proposition 4.7 by noting that $c=\infty$ here and any continuous piecewise affine interpolation is in $\mathcal{H}$.
$\Leftarrow$”: Let $u_{\theta}$ be a solution of (19) with $c=\infty$. We define $u_{\text{FD}}\in\mathbb{R}^{N}$ such that $u_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$. This implies that $E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$.
Assume that $u_{\text{FD}}$ is not a minimizer of $E_{\text{FD}}$, i.e. there is a $\tilde{u}_{\text{FD}}\in\mathbb{R}^{N}$ with $\tilde{u}_{\text{FD}}\neq{u}_{\text{FD}}$ such that $E_{\text{FD}}(\tilde{u}_{\text{FD}})<E_{\text{FD}}(u_{\text{FD}})$. We construct a continuous piecewise affine interpolation $\tilde{u}_{\theta}$ of $\tilde{u}_{\text{FD}}$ on $\mathcal{T}$ such that $\tilde{u}_{\text{FD}}(x_{i})=\tilde{u}_{\theta}(x_{i})$ for $i=1,\ldots,N$. Since $\mathcal{H}$ includes any continuous piecewise affine function on $\mathcal{T}$, we have $\tilde{u}_{\theta}\in\mathcal{H}$. Consequently $E_{\theta}(\tilde{u}_{\theta})=E_{\text{FD}}(\tilde{u}_{\text{FD}})<E_{\text{FD}}(u_{\text{FD}})=E_{\theta}(u_{\theta})$, which is a contradiction to $u_{\theta}$ being a solution of (19) with $c=\infty$ and hence $u_{\text{FD}}$ is indeed a minimizer of $E_{\text{FD}}$. ∎
A direct consequence of proposition 4.8 is that (33) does not have a solution if and only if (19) with $c=\infty$ does not have a solution.
@@
~:=$\widetilde{\mathcal{H}}:=\{h:\Omega\to\mathbb{R}\colon\forall x,y\in\Omega^{h},x\neq y,h(x)\text{ and }h(y)\text{ are independent}\},$
@@
Note that $\widetilde{\mathcal{H}}$ in theorem 4.9 can be a piecewise polynomial space of any polynomial order and hence we could set $\widetilde{\mathcal{H}}=\mathcal{H}$ the set of ReLU-NNs as well. Let us use theorem 4.9 to connect physics-informed neural networks (PINNs) and the Deep Ritz Method to a finite difference method by a simple example. To this end we only need to specify the function $\mathcal{G}$ of theorem 4.9. The equivalence of the respective solutions in the sense of theorem 4.9 follows directly from theorem 4.9.
@@ 5.1 Convergence to analytic solution
$g(x)=\begin{cases}0&\text{if }x\in[a,x_{1}]\\ 1&\text{if }x\in(x_{1},b],\end{cases}$
@@
$u^{*}(x)=\begin{cases}0.2&\text{if }x\in[a,x_{1}]\\ 0.8&\text{if }x\in(x_{1},b]\end{cases}$
yielding $E(u^{*})=0.9$, where $E$ is the functional in (1). To numerically compute a solution of (1) within this setting we consider (19) with $N=1000$, $\nabla^{h}=\nabla_{F}^{h}$ being the forward difference gradient with homogeneous Neumann boundary conditions and $|\cdot|_{\blacktriangleright}=|\cdot|$. The neural network architecture, i.e. the solution space $\mathcal{H}$ is specified as follows: any network consists of 2 hidden layers, whereby the first hidden layer has 64 neurons and the second hidden layer has 128 neurons. We use Adam with learning rate 0.01 and numerically compute for $c\in\{0,0.1,0.5,1,10,100\}$ a solution of (19), whereby the respective numerically computed solution is denote by $u_{c}$. The overall algorithm is terminated after 100001 iterations.
@@
Althought the universal approximation results in theorems 3.1 and 3.2 state that for $d=1$ only one hidden layer is needed to approximate any $BV$-function arbitrarily exact, we choose 2 hidden layers in our experiment. The reasons for this are, that it is not clear how many neurons one needs in the hidden layer and the optimization process seems to find easier a solution when 2 hidden layers are used. Note that the optimizer is stochastic and may get stuck at certain local minima, which are far away from the global solution. Indeed we observe this when only one hidden layer is used. For a similar reason we use 3 hidden layers for $d=2$ to obtain a numerical solution, which we will present next.
@@
||$|\cdot|_{\blacktriangleright}\in\{|\cdot|_{2,\gamma},|\cdot|_{2,1,\gamma},\max\{|\cdot|_{2},\gamma\},|\cdot|_{2,1,\max_{\gamma}},\phi_{2,\gamma}(\cdot),\phi_{2,1,\gamma}(\cdot)\}$
instead of $|\cdot|_{2}$ and $|\cdot|_{2,1}$, which did not work as mentioned above. We set $\nabla^{h}=\nabla_{\text{F}}^{h}$ if we use $|\cdot|_{2,\gamma},\max\{|\cdot|_{2},\gamma\}$ or $\phi_{2,\gamma}(\cdot)$, while we set $\nabla^{h}=\nabla_{\text{FB}}^{h}$ if we use $|\cdot|_{2,1,\gamma},|\cdot|_{2,1,\max_{\gamma}}$ or $\phi_{2,1,\gamma}(\cdot)$. In all these cases we choose $\gamma=10^{-10}$. Moreover, in $\nabla^{h}$ homogeneous Dirichlet boundary conditions are incorporated to mimic $\Omega=\mathbb{R}^{2}$ and that the solution is 0 outside the disk, while numerically the domain $[0,1]^{2}$ is chosen. The neural network architecture, i.e. the solution space $\mathcal{H}$ is specified as follows: the network consists of 3 hidden layers while each layer has 128 neurons. Further we use Adam with learning rate 0.001 and the overall algorithm is terminated after 300001 iterations. With this specifications we compute solutions $u_{c}$ for pairs $(N,c)\in\{(33\times 33,0),(65\times 65,1),(129\times 129,10),(257\times 257,100),(513\times 513,1000),(1025\times 1025,10000)\}$ to investigate numerically the $\Gamma$-convergence. In figure 5 we depict the progress of the energy with increasing $(N,c)$ and in figure 6 the difference between the computed approximation and the exact solution projected on the respective grid is presented. We clearly see that with increasing $N$ and $c$ in all considered cases the energy approaches $E(u^{*})$, while the numerical approximations go towards the exact minimizer. The respective obtained numerical solutions are presented in figure 7. We observe that for all considered types of TV-smoothings, i.e. Huber-regularization, $\gamma$-lifting and $\max$-$\gamma$-lifting, $\operatorname{TV}_{2}^{h}$ and $\operatorname{TV}_{2,1}^{h}$ with $(N,c)=(1025\times 1025,10000)$ obtain a rather close approximation of the analytic exact solution. Actually all these approximated discrete total variations seem to behave very similarly. Note that the convergence result from theorems 4.1 and 4.4 only cover $\operatorname{TV}_{2,1}^{h}$ and its smoothed versions. However, our numerical experiments suggest that also for $\operatorname{TV}_{2}^{h}$ and its smoothed version a similar $\Gamma$-convergence result may hold.

### v2

@@ 2. $d=2$
    : Here we will consider two different discretised gradient operators. The standard forward difference discretized gradient operator denoted by $\nabla^{h}_{\text{F}}$ and the above suggested forward-backward discretisation $\nabla^{h}_{\text{FB}}$. For both choices we incorporate in a standard way respective boundary conditions. The two-dimensional discrete total variations read then as
@@ 4.4 Identifying and solving a numerical issue
We recall that $\mathcal{H}^{M}\subset W^{1,1}(\Omega)$ and hence for any $u\in\mathcal{H}^{M}$ we have $\operatorname{TV}(u)=\int_{\Omega}|\nabla u|\text{d}x$. We analyze now the $\Gamma$-convergence of (12) with $\operatorname{TV}(u)=\int_{\Omega}\varphi_{\gamma}(\nabla u)\text{d}x$, where $\varphi_{\gamma}(\cdot)\in\left\{|\cdot|_{2,\gamma},\max\{|\cdot|_{2},\gamma\},\phi_{2,\gamma}(\cdot)\right\}$ and how their minimizers behave for $\gamma\to 0$.
@@ 4.4.1 $\Gamma$-convergence of total variation approximations
Let $\varphi_{\gamma}(\cdot)\in\left\{|\cdot|_{2,\gamma},\max\{|\cdot|_{2},\gamma\},\phi_{2,\gamma}(\cdot)\right\}$, $(\gamma_{m})_{m\in\mathbb{N}}>0$ be monotonically decreasing sequences with $\lim_{m\to\infty}\gamma_{m}=0$ and $A\in\{\mathcal{A}_{h}\cap\mathcal{H}_{c}^{M},\mathcal{H}_{c}^{M}\}$, where $c\geq 0$, $M\in\mathbb{N}$, $h>0$, and $\mathcal{A}_{h}$ defined as in (20). Further we define the functionals $\mathcal{E}_{m}^{c},\mathcal{E}^{c}:L^{1}(\Omega)\to\overline{\mathbb{R}}$ as
@@
where $A_{n}\in\{\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}},\mathcal{H}_{c_{n}}^{M_{n}}\}$. Then there exist sequences $(c_{n})_{n}\subset\mathbb{R}$ and $(M_{n})_{n}\subset\mathbb{N}$ such that the sequence $(\mathcal{E}_{n,m})_{n,m}$ $\Gamma$-converge to $\mathcal{E}$ with respect to the $L^{1}(\Omega)$ topology.
@@
$\begin{split}\min_{u_{\theta}\in\mathcal{H}_{c}^{M}}\frac{\alpha_{1}}{N}\sum_{x\in\Omega^{h}}|(T^{h}u_{\theta}(x)-g(x)|&+\frac{\alpha_{2}}{N}\sum_{x\in\Omega^{h}}|(T^{h}u_{\theta}(x)-g(x)|^{2}\\ &+\frac{\lambda}{N}\sum_{x\in\Omega^{h}}\widetilde{\varphi}_{\gamma}(\nabla_{\text{FB}}^{h}u(x))\end{split}$ (32)
@@ 4.5 Connection to finite difference discretisation
where $g_{\text{FD}}\in\mathbb{R}^{N}$ is a finite difference discretisation of $g$ such that $g_{\text{FD}}(x)=g(x)$ for all $x\in\Omega^{h}$. In particular, if $\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}}=\mathcal{A}_{h_{n}}$ in (24), then (24) is also the integral counterpart of (33). Note that here as in (19) $\nabla^{h}$ and $|\cdot|_{\blacktriangleright}$ might be any suitable discrete gradient and discrete norm. For example, one may choose $\nabla_{\text{F}}^{h}$ and $|\cdot|_{2}$ or $\nabla_{\text{FB}}^{h}$ and $|\cdot|_{2,1}$. Note that if $u_{\text{FD}}(x)=u_{\theta}(x)$ for all $x\in\Omega^{h}$, then $E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$. This observation leads to the following result:
@@
Given the set of points $\{x_{i}\}_{i=1}^{N}=\Omega^{h}\subset\Omega$. If $u_{\text{FD}}\in\mathbb{R}^{N}$ is a solution of (33), then there exists a constant $c\geq 0$ and an $M\in\mathbb{N}$ such that $u_{\theta}\in\mathcal{H}_{c}^{M}$ is a solution of (19), whereby ${u}_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$.
@@
We construct a mesh $\mathcal{T}$ of simplices with vertices $\{x_{i}\}_{i=1}^{N}$ and let $u_{\text{FD}}$ be a solution of (33). Then there exists a $c\geq 0$ and an $M\in\mathbb{N}$ such that $u_{\theta}\in\mathcal{H}_{c}^{M}$ is its continuous piecewise affine interpolation on $\mathcal{T}$, i.e. $u_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$. Hence we have $E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$.
Assume that $u_{\theta}$ is not a minimizer of $E_{\theta}$, i.e. there is a $\tilde{u}_{\theta}\in\mathcal{H}_{c}^{M}$ with $\tilde{u}_{\theta}\neq{u}_{\theta}$ such that $E_{\theta}(\tilde{u}_{\theta})<E_{\theta}(u_{\theta})$. We define $\tilde{u}_{\text{FD}}\in\mathbb{R}^{N}$ as $\tilde{u}_{\text{FD}}(x_{i})=\tilde{u}_{\theta}(x_{i})$ for $i=1,\ldots,N$. Hence we obtain $E_{\text{FD}}(\tilde{u}_{\text{FD}})=E_{\theta}(\tilde{u}_{\theta})<E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$, which is a contradiction showing that $u_{\theta}$ is indeed a minimizer of $E_{\theta}$. ∎
Note that for any $c\geq 0$ and $M\in\mathbb{N}$ (19) has a solution, see theorem 3.4, while this is not the case for (33). This is why the other direction in proposition 4.7 does not hold in general. Moreover, although [7, Theorem 2.1] states that any continuous piecewise affine function can be represented by a ReLU-NN, it seems not possible to preset a $c\geq 0$ and $M\in\mathbb{N}$ such that $\mathcal{H}_{c}^{M}$ contains any continuous piecewise affine functions, not even on a given grid. However, if (33) has a solution then there is a $c\geq 0$ and an $M\in\mathbb{N}$, such that the sets of solutions of (33) and (19), respectively, are equivalent in the sense of proposition 4.7.
If $c=\infty$, i.e. the constraint $|\theta|_{\infty}<c$ in (19) is dropped, and $M=\infty$, i.e. the number of neurons in the hidden layers can be arbitrary large and unbounded, then we even have the following statement.
@@
Given the set of points $\{x_{i}\}_{i=1}^{N}=\Omega^{h}\subset\Omega$ and let $c=\infty$ and $M=\infty$ in (19), then $u_{\text{FD}}\in\mathbb{R}^{N}$ is a solution of (33), if and only if $u_{\theta}\in\mathcal{H}^{M}$ is a solution of (19), whereby ${u}_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$.
@@
$\Rightarrow$”: This direction is shown as proposition 4.7 by noting that $c=\infty=M$ here and any continuous piecewise affine interpolation is in $\mathcal{H}^{M}$.
$\Leftarrow$”: Let $u_{\theta}$ be a solution of (19) with $c=\infty$ and $M=\infty$. We define $u_{\text{FD}}\in\mathbb{R}^{N}$ such that $u_{\text{FD}}(x_{i})=u_{\theta}(x_{i})$ for $i=1,\ldots,N$. This implies that $E_{\theta}(u_{\theta})=E_{\text{FD}}(u_{\text{FD}})$.
Assume that $u_{\text{FD}}$ is not a minimizer of $E_{\text{FD}}$, i.e. there is a $\tilde{u}_{\text{FD}}\in\mathbb{R}^{N}$ with $\tilde{u}_{\text{FD}}\neq{u}_{\text{FD}}$ such that $E_{\text{FD}}(\tilde{u}_{\text{FD}})<E_{\text{FD}}(u_{\text{FD}})$. We construct a continuous piecewise affine interpolation $\tilde{u}_{\theta}$ of $\tilde{u}_{\text{FD}}$ on $\mathcal{T}$ such that $\tilde{u}_{\text{FD}}(x_{i})=\tilde{u}_{\theta}(x_{i})$ for $i=1,\ldots,N$. Since $\mathcal{H}^{M}$ includes any continuous piecewise affine function on $\mathcal{T}$, we have $\tilde{u}_{\theta}\in\mathcal{H}^{M}$. Consequently $E_{\theta}(\tilde{u}_{\theta})=E_{\text{FD}}(\tilde{u}_{\text{FD}})<E_{\text{FD}}(u_{\text{FD}})=E_{\theta}(u_{\theta})$, which is a contradiction to $u_{\theta}$ being a solution of (19) with $c=\infty=M$ and hence $u_{\text{FD}}$ is indeed a minimizer of $E_{\text{FD}}$. ∎
A direct consequence of proposition 4.8 is that (33) does not have a solution if and only if (19) with $c,M=\infty$ does not have a solution.
@@
~:=$\widetilde{\mathcal{H}}:=\{u:\Omega\to\mathbb{R}\colon\forall x,y\in\Omega^{h},x\neq y,u(x)\text{ and }u(y)\text{ are independent}\},$
@@
Note that $\widetilde{\mathcal{H}}$ in theorem 4.9 can be a piecewise polynomial space of any polynomial order and hence we could set ~$\widetilde{\mathcal{H}}=\mathcal{H}^{M}$ the set of ReLU-NNs as well. Let us use theorem 4.9 to connect physics-informed neural networks (PINNs) and the Deep Ritz Method to a finite difference method by a simple example. To this end we only need to specify the function $\mathcal{G}$ of theorem 4.9. The equivalence of the respective solutions in the sense of theorem 4.9 follows directly from theorem 4.9.
@@ 5.1 Convergence to analytic solution
$g(x)=\begin{cases}0&\text{if }x\in[a,x_{1}],\\ 1&\text{if }x\in(x_{1},b],\end{cases}$
@@
$u^{*}(x)=\begin{cases}0.2&\text{if }x\in[a,x_{1}],\\ 0.8&\text{if }x\in(x_{1},b],\end{cases}$
yielding $E(u^{*})=0.9$, where $E$ is the functional in (1). To numerically compute a solution of (1) within this setting we consider (19) with $N=1000$, $\nabla^{h}=\nabla_{F}^{h}$ being the forward difference gradient with homogeneous Neumann boundary conditions and $|\cdot|_{\blacktriangleright}=|\cdot|$. The neural network architecture, i.e. the solution space $\mathcal{H}^{M}$ is specified as follows: any network consists of 2 hidden layers, whereby the first hidden layer has 64 neurons and the second hidden layer has 128 neurons. We use Adam with learning rate 0.01 and numerically compute for $c\in\{0,0.1,0.5,1,10,100\}$ a solution of (19), whereby the respective numerically computed solution is denote by $u_{c}$. The overall algorithm is terminated after 100001 iterations.
@@
Althought the universal approximation results in theorems 3.1 and 3.2 state that for $d=1$ only one hidden layer is needed to approximate any $BV$-function arbitrarily exact, we choose 2 hidden layers in our experiment. The reasons for this are, that it is not clear how many neurons one needs in the hidden layer and the optimization process seems to find easier a solution when 2 hidden layers are used. Note that the optimizer is stochastic and may get stuck at certain local minima, which are far away from the global solution. Indeed we observe this when only one hidden layer is used. For a similar reason we use 3 hidden layers in our experiments for $d=2$, which we will present next.
@@
||$|\cdot|_{\blacktriangleright}\in\left\{|\cdot|_{2,\gamma},|\cdot|_{2,1,\gamma},\max\{|\cdot|_{2},\gamma\},|\cdot|_{2,1,\max_{\gamma}},\phi_{2,\gamma}(\cdot),\phi_{2,1,\gamma}(\cdot)\right\}$
instead of $|\cdot|_{2}$ and $|\cdot|_{2,1}$, which did not work as mentioned above. We set $\nabla^{h}=\nabla_{\text{F}}^{h}$ if we use $|\cdot|_{2,\gamma},\max\{|\cdot|_{2},\gamma\}$ or $\phi_{2,\gamma}(\cdot)$, while we set $\nabla^{h}=\nabla_{\text{FB}}^{h}$ if we use $|\cdot|_{2,1,\gamma},|\cdot|_{2,1,\max_{\gamma}}$ or $\phi_{2,1,\gamma}(\cdot)$. In all these cases we choose $\gamma=10^{-10}$. Moreover, in $\nabla^{h}$ homogeneous Dirichlet boundary conditions are incorporated to mimic $\Omega=\mathbb{R}^{2}$ and that the solution is 0 outside the disk, while numerically the domain $[0,1]^{2}$ is chosen. The neural network architecture, i.e. the solution space $\mathcal{H}^{M}$ is specified as follows: the network consists of 3 hidden layers while each layer has 128 neurons. Further we use Adam with learning rate 0.001 and the overall algorithm is terminated after 300001 iterations. With this specifications we compute solutions $u_{c}$ for pairs $(N,c)\in\{(33\times 33,0),(65\times 65,1),(129\times 129,10),(257\times 257,100),(513\times 513,1000),(1025\times 1025,10000)\}$ to investigate numerically the $\Gamma$-convergence. In figure 5 we depict the progress of the energy with increasing $(N,c)$ and in figure 6 the difference between the computed approximation and the exact solution projected on the respective grid is presented. We clearly see that with increasing $N$ and $c$ in all considered cases the energy approaches $E(u^{*})$, while the numerical approximations go towards the exact minimizer. The respective obtained numerical solutions are presented in figure 7. We observe that for all considered types of TV-smoothings, i.e. Huber-regularization, $\gamma$-lifting and $\max$-$\gamma$-lifting, $\operatorname{TV}_{2}^{h}$ and $\operatorname{TV}_{2,1}^{h}$ with $(N,c)=(1025\times 1025,10000)$ obtain a rather close approximation of the analytic exact solution. Actually all these approximated discrete total variations seem to behave very similarly. Note that the convergence results from theorems 4.1 and 4.4 only cover $\operatorname{TV}_{2,1}^{h}$ and its smoothed versions. However, our numerical experiments suggest that also for $\operatorname{TV}_{2}^{h}$ and its smoothed versions a similar $\Gamma$-convergence result may hold.